OpenStudy (anonymous):

Find the slope of the line passing through the points (0, -4) and (-6, 7).

OpenStudy (anonymous):

by definition slope m = (y1-y)/(x1-x)

OpenStudy (anonymous):

Check out this question: http://openstudy.com/users/radq#/users/radq/updates/4dfb73cc0b8bbe4f12e64668

OpenStudy (anonymous):

then: m=(7-(-4))/(-6-0)=-3/6=-1/2

OpenStudy (anonymous):

the answer choices are 11/6, -11/6, -6/11, or 6/11

OpenStudy (anonymous):

11/6

OpenStudy (anonymous):

sorry, -11/6

OpenStudy (amistre64):

subtract one point from the other: (0, -4) - (-6, 7) --------- 6 , -11 ; now slope = y/x = -11/6